OpenStudy (anonymous):

Which function below has the largest slope? f(x) x f(x) -3 -10 -1 -7 1 -4 g(x) g(x) over g(x) equals one half x plus 3 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10. f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

well we dont know the units involved so are we allowed to assume it is just counting units like 1 2 3...

OpenStudy (anonymous):

THis is simple i think i am doing it right but whe have to find m=slope for each function. For f(x) we use the value taple and  m=1.5 --- now g(x) equals one half x plus 3 = 1/2 x + 3 y = m x + b Imidiately we see m=slope = 1/2 --- Lance gains 1 level each time he plays his favorite video game. so that can be interpreted as 1/1 or just 1. m=1 --- so f(x) has highest cuzz it is 1.5 I'd wait for second opinion tho.